# Importance Sampling-Guided Meta-Training for Intelligent Agents in Highly Interactive Environments

### v1

@@ Abstract
Training intelligent agents to navigate highly interactive environments presents significant challenges. While guided meta reinforcement learning (RL) approach that first trains a guiding policy to train the ego agent has proven effective in improving generalizability across various levels of interaction, the state-of-the-art method tends to be overly sensitive to extreme cases, impairing the agents’ performance in the more common scenarios. This study introduces a novel training framework that integrates guided meta RL with importance sampling (IS) to optimize training distributions for navigating highly interactive driving scenarios, such as T-intersections. Unlike traditional methods that may underrepresent critical interactions or overemphasize extreme cases during training, our approach strategically adjusts the training distribution towards more challenging driving behaviors using IS proposal distributions and applies the importance ratio to de-bias the result. By estimating a naturalistic distribution from real-world datasets and employing a mixture model for iterative training refinements, the framework ensures a balanced focus across common and extreme driving scenarios. Experiments conducted with both synthetic dataset and T-intersection scenarios from the InD dataset demonstrate not only accelerated training but also improvement in agent performance under naturalistic conditions, showcasing the efficacy of combining IS with meta RL in training reliable autonomous agents for highly interactive navigation tasks.
@@ I Introduction
Autonomous agents are often tasked with navigating complex, highly interactive environments, such as congested and unsignalized intersections [isele2018navigating, wei2021autonomous, ma2021reinforcement, li2024multi, anon2024multi, li2024interactive]. The direct training of these agents using a naturalistic distribution of driving scenarios is notably inefficient due to the imbalanced frequency of scenarios—common scenarios are overrepresented while critical, interactive scenarios are rare yet essential for training [arief2018accelerated, ding2023survey].
It is an open question of how to best train agents when the distribution associated with their target operation is unknown [schaul2015prioritized, isele2018selective]. Recent developments have highlighted the effectiveness of guided meta reinforcement learning (RL) in training agents to handle a variety of interactive driving situations [lee2023robust]. Here, the core strategy involves training the ego agent to interact with a range of social agents, each governed by uniquely parameterized reward functions. These functions are designed to emulate a broad spectrum of cooperative or adversarial behaviors, thus aiming to mimic and challenge human-like driving patterns to increase the robustness of the ego agent’s policy against naturalistic driving conditions. Nonetheless, this method tends to overemphasize extreme scenarios that might not be as frequent in real-world driving but are sampled disproportionately during training, leading to performance degradation for more common driving conditions [lee2023robust, li2024adaptive].
To address this issue, strategies such as oversampling and undersampling have been discussed in the literature [zhou2020oversampling, hong2023beyond]. Implementing these methods in an online training environment remains challenging. Techniques from variance reduction, such as using control variates for efficient gradient estimation [liu2019taming] and importance sampling (IS) for efficient evaluation sampler have been proposed. The main idea of IS in this scheme is to skew the sampling towards extreme scenarios and then adjust the outcomes with importance ratios for an unbiased result. While IS-based approaches have shown appealing results, these approaches are typically applied in evaluation phases [zhao2016accelerated, o2018scalable, arief2022certifiable]. Whether these approaches could be adapted to enhance training remains largely underexplored. f
@@
This study introduces a framework that employs IS both during training and evaluation to mitigate the challenge of overemphasis on extreme scenarios. The proposed training framework integrates a guided meta RL training approach with IS, optimizing the training distribution to efficiently sample critical interactive scenarios without disproportionately emphasizing these scenarios. The IS-optimized training approach strategically biases sampling towards more intense driving situations using an IS proposal derived through the cross-entropy method [kroese2013cross], then computes the importance ratio based on this proposal and the underlying naturalistic distribution to provide an unbiased reward estimate during training. We validate our approach using real-world data from T-intersection scenarios in the InD dataset [inDdataset], demonstrating that it not only enhances the training process but also maintains a balanced emphasis on both common and critical driving scenarios.
@@
  • We integrate the IS proposal, often used for evaluation, into an optimized training distribution using both cross-entropy and mixture models.
  • We introduce an IS-guided meta RL policy training framework that effectively balances the training for both common and critical driving scenarios, enhancing the adaptability and efficacy of the training process by reflecting real-world driving dynamics accurately.
  • We demonstrate through simulation with probability distributions learned from a real-world dataset that our approach significantly improves the generalizability of the trained policies across diverse driving situations.
These contributions collectively enhance the adaptability and efficacy of autonomous driving agents, paving the way for safer and more reliable autonomous vehicle operations in highly interactive real-world conditions.
@@ II Related Work
This section briefly reviews earlier work in policy learning, IS-based evaluation techniques, and interactive autonomous navigation tasks.
@@ II-A Policy Learning for Autonomous Navigation
A broad spectrum of planning and policy learning frameworks have been proposed, evolving from earlier models such as simple Markov Decision Processes (MDPs) [guan2018markov] to more sophisticated structures like Partially Observable MDPs (POMDPs) [qiao2018pomdp, sunberg2022improving] and deep RL methods [zhu2021survey, li2024interactive]. These frameworks have contributed significantly to the development of autonomous driving policies by enabling the modeling of uncertainty and partial observability, which are crucial in dynamic driving environments. Training diverse policies under these frameworks is achieved through reward randomization techniques [tang2021discovering, mysore2021multi], resulting in policies capable of handling complex scenarios such as navigating through uncontrolled intersections. Additionally, adversarial training methods have been incorporated to further strengthen these policies against uncertainties and adversarial conditions [chen2021adversarial].
Despite these advancements, effectively addressing rare critical scenarios remains a challenge. Techniques such as conditional flows [ding2023causalaf], adaptive scenario-set sampling [du2019finding, zhang2022finding], and bridge sampling [sinha2020neural] have been developed to enhance the sampling of these scenarios. These methods aim to improve the robustness and reliability of autonomous driving policies by ensuring that they can handle a wide range of potential driving situations. However, the complexity of these methods poses significant integration challenges when used during the training phase. There is also a persistent risk that policies may become excessively biased towards rare, extreme cases, potentially compromising the agent’s overall performance in more common driving scenarios [lee2023robust].
@@ II-B IS-based Evaluation Approaches
Importance Sampling (IS) has been used to prioritize sampling in various applications [ichter2018learning, luo2019importance] and has emerged as a crucial technique in evaluating the effectiveness of autonomous agent policies under rare and extreme scenarios. IS achieves this by sampling aggressive scenarios more frequently and applying likelihood ratios to ensure unbiased performance estimations [kroese2013cross, kim2016improving, o2018scalable, corso2020survey]. The aggressive scenario distribution used in IS, referred to as the IS proposal distribution, can be refined through methods like cross-entropy (CE) [kroese2013cross, o2018scalable], dominating point analysis with mixture models [huang2018versatile, arief2022certifiable], and failure-based normalizing flows [gibson2023flow].
While previous studies have shown the potential of addressing rare, extreme scenarios during planning [kim2017heuristics], a significant challenge remains on how to effectively bridge the gap between evaluation techniques and actual policy training. This disconnect often results in trained policies that do not fully align with real-world demands. Effective integration of IS-based evaluation approaches with policy learning methods is crucial to developing robust autonomous driving systems. There is a need for integrated approaches that can translate these evaluation techniques into practical training methodologies, ensuring that trained policies are not only theoretically sound but also practically viable in real-world driving conditions.
@@ II-C Autonomous Agents in Highly Interactive Environments
Recent studies have also explored the use of social dynamics and negotiation strategies to enhance the interaction capabilities of autonomous agents. For instance, game theory can be used to model interactions among multiple agents allowing for the prediction and anticipation of other agents’ actions, leading to more effective decision making processes [bellemare2016unifying, yuan2021deep, li2022efficient, li2023game]. Additionally, incorporating techniques such as imitation learning and behavior cloning has enabled agents to learn from human demonstrations, thereby improving their ability to interact seamlessly with human drivers and pedestrians [hussein2017imitation, wang2021imitation].
In this work, we design a framework that can robustly handle highly interactive and dynamic environments without sacrificing its nominal performance. Existing approaches often struggle with scalability and generalization, particularly when dealing with rare, critical interactions or highly unpredictable scenarios. Our research stands out by proposing a framework that integrates IS methods with meta RL to enhance the ability to navigate complex environments and effectively interact with other agents.
@@ III Framework
In this section, we present the proposed framework that integrates IS in both policy evaluation and training for autonomous driving. The framework uses an optimized IS proposal to enhance both the evaluation and subsequent training efficiency of autonomous driving agents. This dual application of IS is pivotal in generating critical scenarios that are not only essential for robust policy assessment but also beneficial for iterative policy enhancement.
@@ III-B Social and Meta-Policy Training
$\max_{\pi\in\Pi_{\text{social}}}\mathds{E}\left[\sum_{t=0}^{\infty}\gamma^{t}(r^{\text{goal}}(s_{t},a_{\text{social},t})+\beta r^{\text{speed}}(s_{t},a_{\text{social},t}))\right],$ (2)
where $a_{\text{social},t}=\pi_{\text{social},\beta}(s_{t})$, and $r^{\text{goal}}$ and $r^{\text{speed}}$ are the reward components for achieving the goal and maintaining speed, respectively.
To train a diverse set of social behaviors, we employ a meta-policy $\pi_{\text{social},\beta}$ using a two-stage approach. In the first stage, baseline policies $\pi_{\text{social},\beta}$ are trained for discrete preferences within a set $\overline{B}=\{\overline{\beta}^{1},\ldots,\overline{\beta}^{m}\}$. Each baseline policy $\pi^{*}_{\text{social},\overline{\beta}^{i}}$ targets a specific behavioral model. In the second stage, the meta-policy $\pi_{\text{social},\beta}$ is then trained by sampling $\beta$ from a continuous distribution $U(\beta_{\text{min}},\beta_{\text{max}})$, and is regularized to approximate the nearest baseline policy using the regularization loss:
@@
where $\mathds{1}$ is an indicator function and $D_{KL}$ is the Kullback-Leibler divergence. This approach facilitates the synthesis of a diverse meta-policy capable of adapting to various social behaviors. Note that $\beta$ is only needed during training.
@@ III-D Evaluating the Ego Policy
The evaluation of $\pi_{\text{ego}}$ is designed to mirror realistic conditions, specifically focusing on the policy’s effectiveness in managing mission-critical failures such as crashes or delays at intersections. We use the CE method to refine the IS proposal distribution $p_{\text{evaluation}}$, which is aimed at generating highly informative and challenging scenarios for robust evaluation.
@@
We initiate the CE algorithm with a Gaussian distribution $N(\mu_{0},\sigma^{2})$, where $\sigma$ is set to a fixed value. The mean is then iteratively adjusted based on the performance data from the lowest 10th percentile reward of the simulated scenarios, ensuring focus on scenarios that reveal potential weaknesses in the ego policy. In each iteration, values of $\beta$ are sampled from this Gaussian distribution to simulate driving scenarios that critically evaluate $\pi_{\text{ego}}$. This iterative optimization process is repeated until the parameters of the distribution stabilize, indicated by a $\mu$ change less than 0.01 between iterations.
@@
To quantify the effectiveness of the ego policy under realistic conditions, we compute the final evaluation metric as
@@
$\pi_{\text{ego}}=\mathds{E}_{\beta}\left[\mathds{1}\{\text{failure}\mid\pi_{\text{ego}}\}\right]=\mathds{E}\left[\hat{\phi}_{\pi_{\text{ego}}}\right].$ (6)
The IS approach ensures that although the scenarios are generated from a biased distribution $p_{\text{evaluation}}$, the final performance estimate remains unbiased.
@@ III-E Augmenting Training Distribution
To refine the training of the ego policy $\pi_{\text{ego}}$, we integrate a Gaussian Mixture Model (GMM) into the training distribution. This GMM uses parameters derived from the collection of IS proposal distributions generated throughout the evaluation phase. The mean vector of the GMM comprises all the means from the distributions $\{p_{\text{evaluation}}\}$, and the standard deviation vector $\sigma$ values. We assign equal weights to each component of the mixture, represented by $1/k$, where $k$ is the number of ego policy training iterations.
This method efficiently uses the diverse and specific scenarios identified during the evaluation phase to enhance the training environment. In addition, the use of the IS-based reward strategy guarantees that the training process yields an unbiased estimate of the ego policy’s performance under real-world driving conditions. This integration ensures that the modifications made during the training phase lead to genuine improvements in performance. The GMM policy is used as our current $p_{\text{optimized}}$ distribution. The framework is summarized in Algorithm 1.
@@ IV Experiments
The goal of the experiment is to benchmark the performance of the proposed training method (CEIS) against baseline methods (GEP and NEP). In addition, we also compare a generic IS-based method (GIS). In the first experiment, we use a toy example to illustrate how the proposed framework handles synthetic rare events. In the second experiment, we show how the proposed framework is versatile and can be used when the naturalistic distribution is fit from a real-world driving dataset.

### v2

@@ Abstract
Training intelligent agents to navigate highly interactive environments presents significant challenges. While guided meta reinforcement learning (RL) approach that first trains a guiding policy to train the ego agent has proven effective in improving generalizability across scenarios with various levels of interaction, the state-of-the-art method tends to be overly sensitive to extreme cases, impairing the agents’ performance in the more common scenarios. This study introduces a novel training framework that integrates guided meta RL with importance sampling (IS) to optimize training distributions iteratively for navigating highly interactive driving scenarios , such as T-intersections or roundabouts. Unlike traditional methods that may underrepresent critical interactions or overemphasize extreme cases during training, our approach strategically adjusts the training distribution towards more challenging driving behaviors using IS proposal distributions and applies the importance ratio to de-bias the result. By estimating a naturalistic distribution from real-world datasets and employing a mixture model for iterative training refinements, the framework ensures a balanced focus across common and extreme driving scenarios. Experiments conducted with both synthetic and naturalistic datasets demonstrate both accelerated training and performance improvements under highly interactive driving tasks.
@@ I Introduction
Autonomous agents often navigate complex, highly interactive environments, such as congested and unsignalized intersections [isele2018navigating, wei2021autonomous, ma2021reinforcement, li2024multi, anon2024multi, li2024interactive]. The direct training of these agents using a naturalistic distribution of driving scenarios is notably inefficient due to the imbalanced frequency of scenarios—common scenarios are overrepresented while critical, interactive scenarios are rare yet essential for training [arief2018accelerated, ding2023survey].
It is an open question of how to best train agents when the distribution associated with their target operation is unknown [schaul2015prioritized, isele2018selective]. Recent developments have highlighted the effectiveness of guided meta reinforcement learning (RL) in training agents to handle a variety of interactive driving situations [lee2023robust]. Here, the core strategy involves training the ego agent to interact with a range of social agents, each governed by uniquely parameterized reward functions. These functions are designed to emulate a broad spectrum of cooperative or adversarial behaviors, thus aiming to mimic human-like driving patterns to increase the robustness of the ego agent’s policy against naturalistic driving conditions. Nonetheless, this method tends to overemphasize extreme scenarios that might not be as frequent in real-world driving but are sampled disproportionately during training, leading to performance degradation for more common driving conditions [lee2023robust, li2024adaptive].
To address this issue, strategies such as oversampling and undersampling have been discussed in the literature [zhou2020oversampling, hong2023beyond]. Implementing these methods in an online training environment remains challenging. Techniques from variance reduction, such as using importance sampling (IS) for efficient evaluation sampler have been proposed. The main idea of IS in this scheme is to skew the sampling towards extreme scenarios and then adjust the outcomes with importance ratios for an unbiased result. While IS-based approaches have shown appealing results, these approaches are typically applied in evaluation phases [zhao2016accelerated, o2018scalable, arief2022certifiable]. Whether these approaches could be adapted to enhance training remains largely underexplored.
@@
This study introduces a framework that employs IS both during training and evaluation to mitigate the challenge of overemphasis on extreme scenarios. The proposed training framework integrates a guided meta RL training approach with IS, optimizing the training distribution to efficiently sample critical interactive scenarios without disproportionately emphasizing them. The IS-optimized training approach strategically biases sampling towards more aggressive driving situations using an IS proposal derived through the cross-entropy method [kroese2013cross], then computes the importance ratio based on this proposal and the underlying naturalistic distribution to provide an unbiased reward estimate. We validate our approach on T-intersection [inDdataset] and roundabout [rounDdataset] scenarios, demonstrating that it not only enhances the training process but also maintains a balanced emphasis on common and critical cases.
@@
  • We integrate the IS proposal, often used for evaluation, into an optimized training distribution.
  • We introduce an IS-guided meta RL policy training framework that effectively balances the training for both common and critical driving scenarios, enhancing the training process by reflecting naturalistic driving dynamics more accurately.
  • We demonstrate the generalizability of the proposed approach in real-world driving datasets, which achieves significant improvement compared to baselines.
@@ II-A Policy Learning for Autonomous Navigation
A broad spectrum of planning and policy learning frameworks have been proposed, evolving from earlier models such as simple Markov Decision Processes (MDPs) [guan2018markov] to more sophisticated structures like Partially Observable MDPs (POMDPs) [qiao2018pomdp, sunberg2022improving] and deep RL methods [zhu2021survey, li2024interactive]. These frameworks have contributed significantly to the development of autonomous driving policies by enabling the modeling of uncertainty and partial observability, which are crucial in dynamic driving environments. Training diverse policies under these frameworks is achieved through reward randomization techniques [tang2021discovering, mysore2021multi], resulting in policies capable of handling complex scenarios such as navigating through uncontrolled intersections.
Despite these advancements, effectively addressing rare critical scenarios remains a challenge. Techniques such as conditional flows [ding2023causalaf], adaptive scenario-set sampling [du2019finding, zhang2022finding], and bridge sampling [sinha2020neural] have been developed to enhance the sampling process. However, the complexity of these methods poses significant integration challenges when used during the training phase. There is also a persistent risk that policies may become excessively biased towards rare, extreme cases, potentially compromising the agent’s overall performance in more common driving scenarios [lee2023robust].
@@ II-B IS-based Evaluation Approaches
Importance Sampling (IS) has been used to prioritize sampling in various applications [ichter2018learning, luo2019importance] and has emerged as a crucial technique in evaluating the effectiveness of autonomous agent policies under rare and extreme scenarios. IS samples aggressive scenarios more frequently and applies likelihood ratios to ensure unbiased performance estimations [kroese2013cross]. The aggressive scenario distribution used in IS, referred to as the IS proposal distribution, can be refined through methods like cross-entropy (CE) [kroese2013cross], dominating point analysis with mixture models [arief2022certifiable], and failure-based normalizing flows [gibson2023flow].
While previous studies have shown the potential of addressing rare, extreme scenarios during planning [kim2017heuristics], a significant challenge remains on how to effectively bridge the gap between evaluation techniques and actual policy training. This disconnect often results in trained policies that do not fully align with real-world conditions. Effective integration of IS-based evaluation approaches with policy learning methods is crucial to ensure that trained policies are not only theoretically sound but also practical.
@@ II-C Autonomous Agents in Highly Interactive Environments
Recent studies have also explored the use of social dynamics and negotiation strategies to enhance the interaction capabilities of autonomous agents. For instance, game theory can be used to model interactions among multiple agents allowing for the prediction and anticipation of other agents’ actions, leading to more effective decision making processes [bellemare2016unifying, yuan2021deep, li2023game]. Additionally, incorporating techniques such as imitation learning and behavior cloning has enabled agents to learn from human demonstrations, thereby improving their ability to interact seamlessly with human drivers and pedestrians [hussein2017imitation, wang2021imitation].
In this work, we design a framework that can robustly handle highly interactive and dynamic environments without sacrificing its nominal performance. Existing approaches often struggle with generalization, particularly when dealing with rare, critical interactions or highly unpredictable scenarios. Our research stands out by proposing a framework that integrates IS methods with meta RL to enhance the ability to navigate complex environments and effectively interact with other agents.
@@ III Framework
The proposed framework uses an optimized IS proposal to enhance both the evaluation and subsequent training efficiency of autonomous driving agents. This dual application of IS is pivotal in generating critical scenarios that are not only essential for unbiased policy assessment but also improves the policy.
@@ III-B Social and Meta-Policy Training
$\max_{\pi\in\Pi_{\text{social}}}\mathds{E}\left[\sum_{t=0}^{\infty}\gamma^{t}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r^{\text{1}}}(s_{t},a_{\text{social},t})+\beta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r^{\text{2}}}(s_{t},a_{\text{social},t}))\right],$ (2)
where $a_{\text{social},t}=\pi_{\text{social},\beta}(s_{t})$, and and $r^{\text{1}}$ and $r^{\text{2}}$ are the primary and secondary reward components, respectively.
To train a diverse set of social behaviors, we employ a meta-policy $\pi_{\text{social},\beta}$ using a two-stage approach. In the first stage, baseline policies $\pi_{\text{social},\beta}$ are trained for discrete preferences within a set $\overline{B}=\{\overline{\beta}^{1},\ldots,\overline{\beta}^{m}\}$. Each baseline policy $\pi^{*}_{\text{social},\overline{\beta}^{i}}$ targets a specific behavioral model parameterized by $\overline{\beta}_{i}$ and is a neural network that uses a set of optimized internal parameters that maximize its reward. These policies are called the guiding policies in the baseline method [lee2023robust]. The term ‘behavioral model’ here refers to a driving policy that models a behavior induced by rewards defined by Equation 2, which, in more general sense, may consider various primary or secondary goals. In this work, however, we only consider the case for behavioral models $\pi_{\text{social},\beta}$, $\pi_{\text{social},\beta^{\prime}}$ for social vehicles that differ only due to $\beta\neq\beta^{\prime}$, where $\beta,\beta^{\prime}\in B\subseteq\mathds{R}$, but uses the same primary and secondary reward components.
In the second stage, the meta-policy $\pi_{\text{social},\beta}$ is then trained by sampling $\beta$ from a continuous distribution $U(\beta_{\text{min}},\beta_{\text{max}})$, and is regularized to approximate the nearest guiding policy using the regularization loss:
@@
where $\mathds{1}$ is an indicator function, $D_{KL}$ is the Kullback-Leibler divergence, and $d$ is a small threshold. We also use $\beta_{\min}=\min_{\beta\in\overline{B}}\beta$ and $\beta_{\max}=\max_{\beta\in\overline{B}}\beta$. With this setting, the two-stage approach facilitates the synthesis of a diverse meta-policy capable of adapting to various social behaviors contained in $\overline{B}$, allowing the social policy to learn driving behaviors characterized by parameter $\beta$ and anchored at the guiding policies defined by $\overline{B}$. Note that $\beta$ is only needed during training.
@@ III-D Evaluating the Ego Policy
The evaluation of $\pi_{\text{ego}}$ is designed to mirror realistic conditions, specifically focusing on the policy’s effectiveness in dealing with failures such as crashes. We use the CE method iteratively to design the IS proposal distribution $p_{\text{evaluation}}$ optimized for generating challenging scenarios for the corresponding ego policy.
@@
We initiate the CE algorithm with a Gaussian distribution $N(\mu_{0},\sigma^{2})$ with some fixed $\sigma$. $\mu_{0}$ is then iteratively adjusted based on the performance data from the lowest 10th percentile reward of the simulated scenarios, ensuring focus on scenarios that reveal potential weaknesses in the ego policy. In each iteration, a handful of $\beta$ values are sampled from this distribution to simulate driving scenarios that critically evaluate $\pi_{\text{ego}}$. This iterative optimization process is repeated until the parameters of the distribution coverges, indicated by an infinitesimal change in subsequent $\mu$’s.
@@
To quantify the effectiveness of the ego policy, we compute the final evaluation metric as
@@
$\mathds{E}_{\beta}\left[\mathds{1}\{\text{failure}\mid\pi_{\text{ego}}\}\right]=\mathds{E}\left[\hat{\phi}_{\pi_{\text{ego}}}\right].$ (6)
This approach ensures that although the scenarios are generated from a biased distribution $p_{\text{evaluation}}$, the final performance estimate remains unbiased.
@@ III-E Augmenting Training Distribution
To refine the training of the ego policy $\pi_{\text{ego}}$, we use a Gaussian Mixture Model (GMM) for the training distribution. The GMM uses parameters derived from the collection of IS proposal distributions generated throughout the evaluation phase. The mean vector of the GMM comprises all the means from the distributions $p_{\text{evaluation}}$, and the standard deviation vector $\sigma$ values. We assign equal weights to each component of the mixture, represented by $1/k$, where $k$ is the number of ego policy training iterations.
This method efficiently uses the diverse and specific scenarios identified during the evaluation phase to enhance the training environment. In addition, the use of the IS-based reward strategy guarantees that the training process yields an unbiased estimate of the ego policy’s performance under naturalistic driving conditions and ensures that the modifications made during the training phase lead to performance improvements, given that the challenging scenarios sampled by the IS proposal can eventually be learned by the ego policy. Otherwise, the training process will not be able to improve the policy and might diminish the performance. To avoid this issue, in the implementation, we filter out the scenarios that are not learnable by the ego policy after a few iterations and reweight the GMM weights accordingly to explore different regions of the scenario space. The framework is summarized in Algorithm 1.
@@ IV Experiments
We evaluated our iterative training approach (CEIS) against three baseline methods: generalized ego policy (GEP), naturalistic ego policy (NEP), and generalized with importance sampling (GIS). Our experiments assess how well each method handles both synthetic and real-world driving scenarios.
To ensure rigorous evaluation, we implement a careful data partitioning strategy. For the ego vehicle, we maintain separate training and testing datasets. The social vehicles train on the complete dataset, while the ego policy follows a more structured approach. For single-iteration methods, we use the full training set. Our iterative approaches require an additional split of the training data - one portion for policy training and another for intermediate evaluation. This intermediate set helps construct both $p_{\text{evaluation}}$ and $p_{\text{training}}$, which guide the policy updates. By maintaining this strict separation between training and testing data, we ensure the final performance metrics for the ego policy remain unbiased.
For training the social vehicle, we set the primary reward $r^{1}$ to correspond to navigating the intersection or roundabout without collision, and the secondary reward $r^{2}$ to maintain high speed. We use $\overline{B}=\{-1,-0.5,\cdots,2.5,3\}$ for the guiding policies, where lower values of $\beta$ corresponds with more aggressive driving behaviors. For anchoring the social policy to the guiding policies, we set the threshold $d=0.5$. Finally, for statistical robustness, we repeat each experimental configuration 10 times and report both means and standard deviations.